OpenStudy (anonymous):

Simplify (4a^(5/3))^3/2

OpenStudy (anonymous):

8a^(5/2)

OpenStudy (anonymous):

\[(a^m)^n=a^{mn}\] \[(4a^{\frac{5}{3}})^{\frac{3}{2}}=4^{\frac{3}{2}}a^{\frac{5}{3}\frac{3}{2}}=8^{\frac{5}{2}}\]